OpenStudy (anonymous):

If a = -12/5, then 5x + 2y = 6 and 3x - ay = 4 are parallel. please help!! true or false.

OpenStudy (texaschic101):

3x - ay = 4 3x - (-12/5)y = 4 3x + 12/5y = 4 -- multiply by the LCD, which is 5, to get rid of the fraction 15x + 12y = 20 now we have : 15x + 12y = 20 12y = -15x + 20 y = -15/12x + 20/12 y = - 3/4x + 5/3 (slope here is -3/4) 5x + 2y = 6 2y = -5x + 6 y = -5/2x + 3 (slope here is -5/2) since the slopes are not the same, these lines are not parallel

OpenStudy (texaschic101):

wait...I made a mistake...hold on

OpenStudy (texaschic101):

15x + 12y = 20 12y = -15x + 20 y = -15/12x + 20/12 y = - 5/4x + 5/3 it didn't change the answer...NOT PARALLEL because the slope of 1 is -5/4 and the slope of the other is -5/2